OpenStudy (christos):

lim x->inf of sqrt(x - 3) - sqrt(x) How would you solve that limit ?

OpenStudy (shamil98):

l'hopital's rule

OpenStudy (shamil98):

also use the limit subtraction law and evaluate them seperately

ganeshie8 (ganeshie8):

Have you tried to rationalize ?

OpenStudy (anonymous):

Try to make sense of \[\frac{\left(\sqrt{x-3}-\sqrt{x}\right) \left(\sqrt{x-3}+\sqrt{x}\right)}{\sqrt{x-3}+\sqrt{x}}=-\frac {3}{\sqrt{x-3}+\sqrt{x}}\] The rest will be obvious